OpenStudy (anonymous):

If Julie has one apple squared and Max has three apples squared, How many apples will there be?

OpenStudy (anonymous):

1^2=1 and 3^2=9 1+9=10 so ten apples

OpenStudy (anonymous):

Thanks ^^

OpenStudy (anonymous):

no problemo